# Unraveling the Age Estimation Puzzle: Comparative Analysis of Deep Learning Approaches for Facial Age Estimation

### v1

@@ Abstract
Comparing different age estimation methods poses a challenge due to the unreliability of published results, stemming from inconsistencies in the benchmarking process. Previous studies have reported continuous performance improvements over the past decade using specialized methods; however, our findings challenge these claims. We argue that, for age estimation tasks outside of the low-data regime, designing specialized methods is unnecessary, and the standard approach of utilizing cross-entropy loss is sufficient. This paper aims to address the benchmark shortcomings by evaluating state-of-the-art age estimation methods in a unified and comparable setting. We systematically analyze the impact of various factors, including facial alignment, facial coverage, image resolution, image representation, model architecture, and the amount of data on age estimation results. Surprisingly, these factors often exert a more significant influence than the choice of the age estimation method itself. We assess the generalization capability of each method by evaluating the cross-dataset performance for publicly available age estimation datasets. The results emphasize the importance of using consistent data preprocessing practices and establishing standardized benchmarks to ensure reliable and meaningful comparisons. The source code is available at Facial-Age-Benchmark.
@@ 1 Introduction
Over the past decade, age estimation has gained considerable interest in biometrics, law enforcement, and child protection. Numerous papers employ deep learning methods for age estimation from facial images, utilizing diverse approaches. For both researchers and practitioners intending to utilize age estimation, it is crucial to know: (i) what is the current state-of-the-art method (ii) what performance can be expected from the method (iii) in what ways it can be improved. However, due to differences in the data-splitting of datasets and the lack of detailed descriptions of all components of the prediction system, answering these questions using the published results is not possible.
This paper aims at answering these questions. To this end, we evaluate and compare recent age estimation approaches in a unified setting and analyze the impact of factors, such as facial alignment, facial coverage, image resolution, model architecture, or pre-training data, on the result. In our paper, we demonstrate that these factors frequently exert a more pronounced impact than the choice of the age estimation method itself.
@@
  • We benchmark and fairly compare recent deep-learning methods for age estimation from facial images. We concentrate on state-of-the-art methods that adapt a generic architecture by changing its last layer or the loss function to suit the age estimation task. Besides the usual intra-class performance, we also evaluate their cross-dataset generalization.
  • We analyze the influence of different components of the facial recognition pipeline. Namely, the influence of (i) facial alignment and preprocessing (ii) model architecture (iii) loss function (iv) pre-training and highlight the need to keep them fixed to ensure a fair comparison. Surprisingly, we find that the influence of the loss function and the last layer, usually the main component that distinguishes different methods, is negligible, see Fig. 1.
  • We make the entirety of our code public, including the data preparation pipeline and implementation of the methods. The published code is valuable to both practitioners who are interested in a state-of-the-art implementation and researchers who are looking for a reliable benchmark to evaluate their novel methods.
@@
Figure 1: Mean Absolute Error (MAE) $\downarrow$ of age estimation methods on the MORPH dataset [23], as reported in existing literature. Random splitting remains the prevalent data splitting strategy and shows consistent performance improvements over time. Subject-exclusive data splitting, which offers more precise evaluations, is rarely employed. With unified subject-exclusive data splitting, all evaluated methods yield comparable results, failing to achieve the performance gains promised by the random splitting.
2 Related Work
This section explores the different age estimation techniques, that we compare in this paper, and examines the benchmarks currently utilized for evaluating them.
2.1 Age Estimation Methods
Classification
Extended binary classification
Fixed distribution learning
Gao et al. [12] (DLDL) approach the task as multi-class classification. However, they encode the label distribution as a normal distribution centered at the true label. Díaz and Marathe [9] (SORD) approach the task similarly, but encode the label distribution as a double exponential distribution centered at the true label. In a follow up to their work [12], Gao et al. [13] (DLDL-v2) propose to also minimize the difference between (i) the true label $y\in\mathcal{Y}$ (ii) the expectation $\mathbb{E}_{\hat{y}\sim f(x)}[\hat{y}]$ of the model output distribution $f(x)$.
Adaptive distribution learning
An approach emerging in recent years is not to model a specific distribution, such as normal or double exponential distribution, but instead, to constrain the model by some statistical measure or a condition. Pan et al. [22] (Mean-Variance) approach the task as standard multi-class classification, but design a loss function that (i) minimizes the squared difference between the expectation $\mathbb{E}_{\hat{y}\sim f(x)}[\hat{y}]$ and the true label $y\in\mathcal{Y}$ (ii) minimizes the variance $\mathbb{E}_{\bar{y}\sim f(x)}\left[(\bar{y}-\mathbb{E}_{\hat{y}\sim f(x)}\left[\hat{y}\right])^{2}\right]$ of the model output distribution $f(x)$. Similarly, Li et al. [17] (Unimodal) design a loss function (i) which constrains the model to output unimodal distributions (ii) concentrates the output distribution around the true label $y\in\mathcal{Y}$.
2.2 Age Estimation Benchmarks
@@
The aforementioned publications evaluate their methods on a number of datasets, with the MORPH [23] dataset being the most commonly used. However, the evaluation procedures are not unified. For instance, OR-CNN [21] randomly divides the dataset into two parts: 80% for training and 20% for testing. No mention is made of a validation set for model selection. Identical splitting protocols are used in [19, 27, 3, 12, 13, 5], but the specific data splits differ between studies. Since the dataset contains multiple images per person; with many of them captured at the same age; there is a possibility for the same individual to be present in both the training and testing sets. This overlap introduces a potential bias, resulting in overly optimistic evaluation outcomes. The degree of data leakage can vary when using random splitting, resulting in certain data splits being more challenging than others. Consequently, comparing different methods and discerning which method truly stands out as the most effective becomes problematic.
@@
Some publications [22, 26] recognize this bias introduced by the splitting strategy and address it by implementing subject-exclusive splitting. This approach ensures that individuals are not included in both the training and testing sets. Standardized data splits are provided for two public age estimation datasets: the ChaLearn Looking at People Challenge 2016 (CLAP2016) dataset [1] and the Cross-Age Celebrity Dataset (CACD2000) [4]. However, CLAP2016 is relatively small, consisting of fewer than 8000 images. On the other hand, CACD2000 has training annotations that may contain noise, and its authors do not recommend its use for evaluating age estimation. Consequently, comparing methods using only these datasets is not satisfactory. Other popular datasets, AgeDB dataset [20], Asian Face Age Dataset (AFAD) [21], and FG-NET [15] also consist of multiple images per person, requiring subject-exclusive splitting. However, they lack any standardized data splits and as such suffer the same issues as MORPH [23].
Motivated by these findings, we aim to fairly compare the methods listed in Section 2.1. To achieve this, we evaluate their performance on standardized data splits, ensuring a unified setup. In an effort to facilitate future comparisons, we make our data splits publicly available.
3 Experimental Setup
For a fair comparison of multiple methods, the same experimental setup should be used for each of them. To accurately recreate an experiment, it is essential to replicate four key components: (i) preprocessing (ii) model architecture (iii) the decision layer and the loss function (iv) the data . To compare different age estimation approaches, we keep components (i), (ii), and (iv) constant while varying the component (iii). This allows us to isolate the impact of the selected method on performance. We compare the methods based on the achieved Mean Absolute Error (MAE).
In addition to comparing age estimation methods, we aim to assess the impact of the individual components. To accomplish this, we establish a baseline approach and systematically modify each component to gauge their respective influence on the results. In this section, we provide an overview of the components and describe our baseline approach. Where relevant, we also review how they are typically addressed in the literature.
3.1 Preprocessing
Unlike other deep learning tasks, where the entire image is often used as input, age estimation models should only be provided with a specific region of the image that corresponds to the face. A typical data preprocessing pipeline for facial age estimation consists of multiple steps. The initial step is face detection. Facial alignment is then performed to align facial landmarks, such as the eyes or corners of the mouth, rectifying variations in head pose, rotation, and scale. Finally, the region of interest is cropped and interpolated to a standardized size. We describe how we handle these steps in more detail below.
Facial Detection & Alignment
For face detection and facial landmark detection we use the RetinaFace model developed by Deng et al. [8]. Our facial alignment procedure involves the following steps. Firstly, we determine a point that serves as the center of the final bounding box. This point is positioned between the center of the eyes and the center of the mouth corners. Secondly, the size of the bounding box is adjusted based on the distance between the eyes for frontal images and the distance between the eyes and mouth for profile images. Lastly, the bounding box is rotated to ensure horizontal alignment of the eyes. To provide a concise explanation, we refrain from presenting an extensive description of the alignment procedure. We encourage interested readers to refer to the accompanying implementation for more details.
While facial alignment defines the positioning, orientation, and scale of facial landmarks, the extent to which the face is visible in an image also needs to be specified. We refer to this notion as facial coverage. It measures how much of the face is shown in an image and can range from minimal coverage, where only the eyes and mouth are visible, to complete coverage, where the entire head is visible. Determining the optimal compromise between complete facial coverage and minimal coverage is not immediately clear. Complete facial coverage provides a comprehensive view of the face, allowing age estimation algorithms to consider a broader range of facial cues. On the other hand, partial coverage may help reduce overfitting by eliminating irrelevant facial cues and features with high variance. For a visual demonstration of various facial coverage levels, refer to Fig. 3. Surprisingly, the concept of facial coverage has received limited attention in age estimation literature. Consequently, the extent of facial coverage utilized in previous studies can only be inferred from the images presented in those works. For instance, Berg et al. [2] seemingly employ minimal coverage, showing slightly more than just the mouth and eyes. The majority of other works [21, 26, 12, 13, 3, 17] tend to adopt partial coverage, where a significant portion of the face, including the chin and forehead, is visible, but not the entire head. In the works of Pan et al. [22], Rothe et al. [25], and Zhang et al. [27], the entire head is shown. In our baseline setup, we have opted for complete facial coverage, where the images encompass the entire head. We further study the effects of facial coverage on model performance in Sec. 4.4.
Image Resolution
After applying the facial detection and alignment pipeline, a suitable square bounding box is chosen to match the desired facial coverage. Subsequently, we extract the bounding box and resize the resulting image to the desired size. In our baseline setup, we use a model input size of $256\times 256$ pixels. However, in practice, we extract a slightly larger bounding box (enlarged by an additional 10%) to facilitate data augmentation without introducing padded pixels. We further study the effects of input resolution on model performance in Sec. 4.4.
Representation
Lastly, we normalize the pixel values of the images. To this end, we subtract the mean and divide by the standard deviation of the color channels computed over ImageNet [7]. Recently, Lin et al. [18] proposed to also transform the image to a different coordinate system and the approach has demonstrated impressive results in the semantic segmentation of faces. We do not utilize any such transformation in our baseline setup, however, we investigate the potential benefits of this transformation in Sec. 4.4.

### v3

# A Call to Reflect on Evaluation Practices for Age Estimation: Comparative Analysis of the State-of-the-Art and a Unified Benchmark

@@ Abstract
Comparing different age estimation methods poses a challenge due to the unreliability of published results stemming from inconsistencies in the benchmarking process. Previous studies have reported continuous performance improvements over the past decade using specialized methods; however, our findings challenge these claims. This paper identifies two trivial, yet persistent issues with the currently used evaluation protocol and describes how to resolve them. We offer an extensive comparative analysis for state-of-the-art facial age estimation methods. Surprisingly, we find that the performance differences between the methods are negligible compared to the effect of other factors, such as facial alignment, facial coverage, image resolution, model architecture, or the amount of data used for pretraining. We use the gained insights to propose using FaRL as the backbone model and demonstrate its effectiveness on all public datasets. We make the source code and exact data splits public on GitHub and in the supplementary material.
@@ 1 Introduction
Age estimation has received significant interest in recent years. However, a closer examination of the evaluation process reveals two underlying issues. First, no standardized data splits are defined for most public datasets, and the used splits are rarely made public, making the results irreproducible. Second, methods often modify multiple components of the age estimation system, making it unclear which modification is responsible for the performance gains.
This paper aims to critically analyze the evaluation practices in age estimation research, highlight the issues, and appeal to the community to follow good evaluation practices to resolve them. We benchmark and fairly compare recent deep-learning methods for age estimation from facial images. We focus on state-of-the-art methods that adapt a generic architecture by changing its last layer or the loss function to suit the age estimation task. Although this may appear restrictive, it is essential to note that most of the methods proposed in the field fall into this category ($\approx 70\%$). By comparing methods that modify only a small part of the network, we aim to ensure a fair evaluation, as the remaining setup can be kept identical. Besides the usual intra-class performance, we also evaluate their cross-dataset generalization, which has been neglected in the age prediction literature so far. Surprisingly, we find that the influence of the loss function and the decision layer on the results, usually the primary component that distinguishes different methods, is negligible compared to other factors.
@@
  • We show that existing evaluation practices in age estimation do not provide consistent results. This leads to obstacles for researchers aiming to advance prior work and for practitioners striving to pinpoint the most effective approach for their application.
  • We define a proper evaluation protocol, offer an extensive comparative analysis for state-of-the-art facial age estimation methods, and publish our code.
  • We show that the performance difference caused by using a different decision layer or training loss is significantly smaller than that caused by other parts of the prediction pipeline.
  • We identify that the amount of data used for pre-training is the most influential factor and use the observation to propose using FaRL [31] as the backbone architecture. We demonstrate its effectiveness on public datasets.
@@
Figure 1: Mean Absolute Error (MAE) $\downarrow$ of age estimation methods on the MORPH dataset, as reported in the existing literature and measured by us, viewed over time. Random splitting remains the prevalent data splitting strategy. The consistent performance improvements over time are attributed in the literature to specialized loss functions for age estimation. Subject-exclusive (identity-disjoint) data splitting is rarely employed. With unified subject-exclusive data splitting and all factors except the loss function fixed, all evaluated methods yield comparable results, failing to achieve the performance gains promised by the random splitting.
2 Issues with Current Evaluation Practices
2.1 Data Splits
@@
Publications focused on age estimation evaluate their methods on several datasets [21, 1, 4, 24, 30, 26, 22]. The most commonly used of these is the MORPH [24] dataset. However, the evaluation procedures between the publications are not unified. For instance, OR-CNN [22] randomly divides the dataset into two parts: 80% for training and 20% for testing. No mention is made of a validation set for model selection. Random splitting (RS) protocol is also used in [20, 29, 3, 13, 14, 5], but the specific data splits differ between studies as they are rarely made public. Since the dataset contains multiple images per person (many captured at the same age), the same individual can be present in both the training and testing sets. This overlap introduces a bias, resulting in overly optimistic evaluation outcomes. The degree of data leakage can vary when using random splitting, making certain data splits more challenging than others. Further, this fundamentally changes the entire setup; rarely will one want to deploy the age estimation system on the people present in the training data. Consequently, comparison of different methods and discerning which method stands out as the most effective based on the published results becomes problematic.
Only some publications [23, 28] recognize this bias introduced by the splitting strategy and address it by implementing subject-exclusive (SE) [23] splitting. This approach ensures that all images of an individual are exclusively in the {enumerate*}[label=(),before= , itemjoin=, , itemjoin*=, or ]
training
validation
testing part. The terminology here is not fully established. One might encounter either identity-disjoint or person-disjoint instead of subject-exclusive in the literature.
To assess how prevalent random splitting (RS) on the MORPH dataset truly is, we conducted a survey of all age estimation papers presented at the CVPR and ICCV since 2013. We found 16 papers focused on age estimation, of which nine use RS, two use SE, five use specialized splits, and three do not utilize MORPH. We further surveyed other research conferences and journals, namely: IJCAI, BMVC, ACCV, IEEE TIP, Pattern Recognit. Lett., Pattern Anal. Appl., and find eight influential age estimation papers that use MORPH. Of those, seven use RS, and one uses a specialized split. By specialized splits, we are referring to non-standard strategies such as ethnically balanced partitions.
Altogether, we discover that only $\approx 10\%$ of papers that utilize MORPH use the SE protocol. This finding is concerning, as MORPH [24] is the most popular dataset used to compare age estimation approaches. Other datasets do not provide a reliable benchmark either, as standardized data splits are provided only for two public age estimation datasets: {enumerate*}[label=(),before= , itemjoin=, , itemjoin*=, and ]
the ChaLearn Looking at People Challenge 2016 (CLAP2016) dataset [1], which is relatively small, consisting of fewer than 8000 images
@@
the Cross-Age Celebrity Dataset (CACD2000) [4], which has noisy training annotations and is not intended for age estimation. Comparing methods using only these datasets is, therefore, not satisfactory either. Other popular datasets, AgeDB dataset [21] and Asian Face Age Dataset (AFAD) [22], also consist of multiple images per person, requiring SE splitting. However, they lack any data splits accepted by the community and often are used with the RS protocol. As such, they suffer from the same issues as MORPH [24].
2.2 Pipeline Ablation
To fairly compare multiple methods, an identical experimental setup should be used for each of them. The current state-of-the-art age estimation approaches adhere to a common framework encompassing: {enumerate*}[label=(),before= , itemjoin=, , itemjoin*=, and ]
data collection
data preprocessing
model design, including the decision layer and the loss function
training and evaluation. Most novel approaches introduce distinct changes to the component (iii); namely they design a specialized loss function to exploit the ordinal nature of age. However, they frequently alter multiple components of the framework simultaneously, complicating the attribution of performance improvements to the claimed modifications.
To compare different loss functions, e.g., [22, 3, 23, 18, 17, 13, 14, 10], the other components of the framework should be kept constant, allowing us to isolate the impact of the selected method on the performance. This is trivial, yet the age estimation community mostly ignores it. Further, many publications hand-wave the other components and do not precisely specify them, making future comparisons meaningless. It’s important to question whether the reported enhancement in a research paper truly stems from the novel loss function it proposes or if it could be attributed to a different modification. We strongly advocate that each component be addressed in isolation and that the experimental setup be precisely described.
Over the past decade, numerous novel age estimation methods have been introduced, promising continuous performance improvements every year. However, motivated by these findings, we raise the question: how reliable are the published age estimation results? In Sec. 3 we aim to establish a proper evaluation protocol and use it in Sec. 4 to compare the methods [25, 13, 14, 23, 18, 10, 22] reliably. Figure 1 illustrates the contrast between the performance of state-of-the-art methods as reported in their respective studies and the outcomes as measured by our implementation.
3 Evaluation Protocol
We identified two trivial yet persistent issues that prevent a reliable comparison of age estimation methods. In this section, we address the initial challenge concerning consistent data partitioning. We provide clear guidelines for the evaluation protocol to ensure replicable and fair assessments. Specifically, the protocol should establish a reproducible approach for defining the data used in both (i) training and (ii) performance evaluation. When specifying the training data, one needs to state whether the training dataset is the sole source of information, or if the model was pretrained with additional data. Additionally, the evaluation can be subdivided based on the data used for model evaluation into intra-dataset, and cross-dataset results. We describe how to evaluate models in these settings below.
Intra-dataset performance
To evaluate intra-dataset performance, a single dataset is used for both training and evaluation of the age estimation system. In this case one should {enumerate*}[label=(),before= , itemjoin=, , itemjoin*=, then ]
randomly split the dataset into subject-exclusive training, validation, and test set  [^1] The generated training, validation and test sets will usualy be a partition of the dataset, however, in any case their intersection must be empty.
train the model on the training set
measure the model’s performance on the validation set
possibly revert back to step (ii) and train the model again
evaluate the model’s performance on the test set
publish the results on the test set along with a detailed description of the system components and the data used. If the dataset consists of limited number of examples, it is possible to create multiple splits of the data into training, validation and test set through step (i). Following this, steps (ii) through (v) are iterated $n$ times, where $n$ is the number of generated splits. It is advisable to present the average test performance along with its standard deviation when reporting the results.
Cross-dataset performance
To evaluate cross-dataset performance, the data split in step (i) of the aforementioned evaluation process is generated from a collection of multiple datasets, ensuring that the complete chosen dataset must be employed entirely for evaluation, effectively constituting the designated test set. The remaning steps of the evaluation procedute remain unaltered.
Regardless of the scenario, whether it is intra-dataset or cross-dataset, each system needs to be evaluated against the test data only once, and the results published. All prior model development and hyperparameter tuning must be based solely on the results on the validation set. Furthermore, it should be indicated whether the training data are the only source of information used for training, or whether the model was pretrained with additional data. In the letter scenario, a detailed description of the additional data and their utilization should also be provided.
4 Comparative Method Analysis
This section applies the evaluation protocol to compare state-of-the-art age estimation methods. We maintain a consistent preprocessing procedure, model architecture, and dataset while selectively altering the decision layer and loss function to incorporate modifications proposed in prominent works such as [25, 13, 14, 23, 18, 10, 22].
4.1 Methodology
Datasets
We evaluate the methods using 7 datasets: AgeDB [21], AFAD [22], CACD2000 [4], CLAP2016 [1], FG-NET [16], MORPH [24], and UTKFace [30]. We also use the IMDB-WIKI dataset [26] for pre-training with clean labels from Franc and Čech [12].
Data Splits
For the CLAP2016 and CACD2000 datasets, we use the single data split provided by the dataset authors. For the remaining datasets, we create five subject-exclusive (SE) data splits. To generate the split, we partition the dataset such that $60\%$ of the dataset is used for training, $20\%$ for model selection (validation), and $20\%$ for evaluating the model performance (test). Additionally, we ensure that each partition has the same age distribution. Due to its small size, we only use FG-NET for evaluation. We make our data splits and code public at Facial-Age-Benchmark [^2] https://github.com/paplhjak/Facial-Age-Estimation-Benchmark.
Model Architecture & Weight Initialization
We use ResNet-50 [15] as the backbone architecture. We always start the training of the methods from the same initialization. We run the experiments with {enumerate*}[label=(),before= , itemjoin=, , itemjoin*=, and ]
random initialization
weights pre-trained on ImageNet (TorchVision’s IMAGENET1K_V2)
weights pre-trained on ImageNet and then further trained on IMDB-WIKI for age estimation with cross-entropy. After the pre-training, the last layer of the model is replaced with a layer specific to the desired method. The models are then fine-tuned on the downstream dataset. It is important to note that for the baseline cross-entropy, we also replace the final layer before fine-tuning. This ensures that the experimental setup remains identical to that of the other methods.
Training Details
We utilize the Adam optimizer with the parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$. For pre-training on the IMDB-WIKI dataset, we set the learning rate to ${\alpha=10^{-3}}$ and train the model for a total of 100 epochs. For fine-tuning on the remaining datasets we reduce the learning rate to ${\alpha=10^{-4}}$ and train the model for 50 epochs. We use a batch size of 100. The best model is selected based on the MAE metric computed on the validation set. We utilize two data augmentations during training, {enumerate*}[label=(),before= , itemjoin=, , itemjoin*=, and ]
horizontal mirroring
cropping out an 80% to 100% portion of the bounding box and resizing it to the model input shape. We do not tune the hyperparameters of the methods [22, 13, 14, 10, 23, 18] on the validation set. We apply them in the original configurations. We argue that if any of the loss functions is a significant improvement over the baseline, we should observe a performance improvement across a broad range of hyperparameters and preprocessing pipelines. We consider our training parameters to be reasonable and to provide a comparison of the methods as if employed out-of-the-box.
Preprocessing
We use the RetinaFace model developed by Deng et al. [8] for face detection and facial landmark detection. We use complete facial coverage, i.e., the images encompass the entire head. We resize the images to a resolution of $256\times 256$ pixels and normalize the pixel values of the images. To this end, we subtract the mean and divide by the standard deviation of colors on ImageNet [7].
Metrics
We use the Mean Absolute Error (MAE) calculated on the test data as the performance measure. To determine whether any method is consistently better than others, we employ the Friedman test and the Nemenyi critical difference test (FN test) as described by Demšar [6]. The main statistic used in the test is the average ranking ($1$ is best) of a method computed on multiple datasets. Differences in the average ranking are then used to decide whether a method is significantly better than others or whether the improvement is due to randomness (the null hypothesis). We use a common significance level (p-value) of $\alpha=5\%$.
4.2 Results
Intra-Dataset Performance
The intra-dataset results can be seen in Tab. 6, highlighted with a grey background. When starting from random initialization, training with the Unimodal loss [18] tends to be unstable. Excluding the Unimodal loss [18] from the evaluation, we apply the FN test. The results indicate that three methods: OR-CNN [22], DLDL [13], and the Mean-Variance loss [23], demonstrate a significant performance improvement over the baseline cross-entropy. With limited data availability, when pre-training is not possible, it is advisable to utilize one of the aforementioned methods.
With pre-training, either on ImageNet or IMDB-WIKI, none of the methods is significantly better than the cross-entropy. In other words, we do not observe any systematic improvement by deviating from the standard approach.
Cross-Dataset Generalization
Cross-dataset results, shown in Tab. 6 with white background, were obtained by evaluating the performance of models on datasets that were not used for their training. The cross-dataset performance is unsurprisingly significantly worse than the intra-dataset performance for all of the methods. Using the FN test, we conclude that there is no significant difference in generalization capability between any of the methods [22, 13, 14, 10, 23, 18] and the cross-entropy, regardless of pre-training. In other words, though the loss functions may reduce overfitting, they do not help in the presence of covariate shift.
None of the methods perform well when evaluated on a different dataset than the one they were trained on. The best cross-dataset results are achieved by training on either UTKFace or CLAP2016. The worst performance across databases is observed when models are trained on AFAD or MORPH. This discrepancy can be attributed to UTKFace and CLAP2016 having a broader range of images, which allows them to generalize effectively to other datasets. Conversely, the limited diversity in MORPH or AFAD datasets, such as AFAD mainly comprising images of people of Asian ethnicity and around $80\%$ of MORPH being composed of individuals of African American ethnicity, contributes to the poor knowledge transfer. The significant decrease in the performance of models trained on the MORPH dataset when applied to other age estimation datasets underscores the importance of not relying solely on the MORPH dataset as the benchmark for age estimation. To ensure a reliable evaluation of different methods, it is crucial to incorporate results from alternative datasets as well.
5 Component Analysis
In this section, we analyze the influence of the backbone architecture and the data preparation pipeline on model performance. We show that changes to these components can have a much more significant impact on the final performance than the choice of a loss function. When altering a component, we maintain all other components at their defaults, presented as the Cross-Entropy approach in Sec. 4. We use the gained insight to propose a strong baseline age estimation model using the FaRL [31] backbone.